OpenStudy (anonymous):

what is equivalent 65/56 as a fraction without a whole number? There is no answer choices

OpenStudy (anonymous):

i will give a medal

OpenStudy (anonymous):

hold on im trying to solve it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1.160714 thats i can figure out. are you doing ratios? or what because i dont understand the qestion

OpenStudy (anonymous):

fractions

OpenStudy (anonymous):

hmm let me think

OpenStudy (anonymous):

is this on a worksheet? in a book? or online?

OpenStudy (anonymous):

online

OpenStudy (anonymous):

can you send me the link to the web page?

OpenStudy (anonymous):

the real question is ''Sue has 7/8 of the book left to read,if she reads another 2/7 of the book how much is left. A)33/56 B)1/5 C)5/8 D)23/40

OpenStudy (anonymous):

ok give me a min

OpenStudy (anonymous):

65/56 = 1.160714 33/56 = 0.589286 1/5 = 0.2 5/8 = 0.625 23/40 = 0.575 can you round?

OpenStudy (anonymous):

yeah why

OpenStudy (anonymous):

ok i put all of the frations in dec. form i posted it cant you round one of them and see which one it would be?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

ok and sorry i couldnt be much more help

OpenStudy (anonymous):

did it work

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

it was 33/56